OpenStudy (anonymous):

use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2+2x-10

OpenStudy (sasogeek):

what is the quadratic formula?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

ax^2+bx+c=0 ?

OpenStudy (sasogeek):

nope :P well the quadratic formula states that \(\huge x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \)

OpenStudy (sasogeek):

can you continue now that you have the formula?

OpenStudy (anonymous):

no

OpenStudy (sasogeek):

ok so if you have \(\huge ax^2+bx+c=0 \) substitute the values of a b and c into the formula and solve it.

OpenStudy (sasogeek):

are we okay now?

OpenStudy (anonymous):

To find x-intercepts put y =0 in y=x^2+2x-10

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nw it is a quad eq....use Quadratic Formula..

OpenStudy (anonymous):

ok how do I do that ?

OpenStudy (sasogeek):

you have an equation similar to ax^2+bx+c=0 substitute the values into the quadratic formula i gave you, and solve it algebraically

OpenStudy (anonymous):

I don not know how to do that ???????

OpenStudy (sasogeek):

a=1 b=2 c=-10 replace that into the equation that into the formula

OpenStudy (sasogeek):

how are we doing?

OpenStudy (anonymous):

x^2+2x+-10=0

OpenStudy (sasogeek):

what about it?

OpenStudy (anonymous):

so now what do I do w/ it ?

OpenStudy (sasogeek):

i told you what a,b, and c are in that equation. a=1 b=2 c=-10 replace the values of a,b and c into the quadratic formula i gave you and solve it to find x

OpenStudy (sasogeek):

which part of that don't you understand? :/